The incidence of retinitis pigmentosa in Britain is about 1:2500. What is the approximate allele frequency in Britain's population?
Answer #1

1/2500

q2=1/2500

q=1/50

q=0.02=allele frequency for retina pigmentosa allele

Unaffected people frequency p=1-q

=1-0.02

=0.98